OpenStudy (anonymous):

What is the volume of 15 g of gaseous 02 at a pressure of 1.15 atm and a temperature of 40˚C? A:1.34 L B:20.95 L C:10.48 L

OpenStudy (aaronq):

use the ideal gas law, PV=nRT rearrange for V, volume.

OpenStudy (jfraser):

be sure to change the grams of O2 into moles, first!

OpenStudy (anonymous):

I DIDNT GET ANY OF THOSE ANSWERES :(

OpenStudy (jfraser):

the ideal gas law is for moles of gas, but you're given grams. COnvert the grams of O2 into moles of O2 first, then plug into the IGL\[P*V = n*R*T\]so\[V = \frac {n*R*T}{P}\]\[moles_{O2} = 15g O_2 *\frac{1mol}{32gO_2}\] take that answer in moles and plug into the IGL with the rest of the numbers

OpenStudy (anonymous):

OK WOULD IT BE B?

OpenStudy (jfraser):

not what i got, sorry

OpenStudy (anonymous):

AWWWWW HOLD UP....

OpenStudy (jfraser):

what did you use for the molar mass of O2?

OpenStudy (anonymous):

OK ITS C :)

OpenStudy (jfraser):

yep

OpenStudy (anonymous):

YAY ^_^ THANKX

OpenStudy (anonymous):

You have a gas sample that is currently at an unknown volume. You need the gas to expand to fill a balloon that has a capacity of 50 L. You know its pressure is 1.05 atm and that you have 1.90 mol of the gas. What temperature does the gas need to be at to fill the balloon? A:337˚C B:337 K C:64 K DO U GUYS KNOW THIS ONE BY ANYCHANCE?

OpenStudy (jfraser):

you asked this in physics already, check my post

OpenStudy (anonymous):

OK BUT I STILL NEED HELP

OpenStudy (jfraser):

rearrange the IGL to solve for T, plug and solve

OpenStudy (anonymous):

ITS B R IGHT :)

OpenStudy (jfraser):

what are the proper unts of temperature for gas laws? A and B are the same number, just not the same UNIT

OpenStudy (anonymous):

SO IT IS B?

OpenStudy (jfraser):

it's B